# THE RESTRICTED PICARD FUNCTOR

### v2

@@ 5 Almost-representability of the restricted Picard functor
In this section we prove the almost-representability of the restricted Picard functor for regular $K$-schemes of finite type. As a dévissage step, we show that, in proving Theorem 1.3, we are free to restrict to an open subscheme.
@@
Let $j\colon U\rightarrow X$ denote the inclusion, and consider the restriction map $j^{*}\colon\operatorname{Pic}^{+}_{X/K}\rightarrow\operatorname{Pic}^{+}_{U/K}$. For every smooth $K$-scheme $T$, $X\times T$ is regular, hence $\operatorname{Pic}(X\times T)\twoheadrightarrow\operatorname{Pic}(U\times T)$. It follows that $j^{*}$ is a surjection of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$. We claim that $\ker(j^{*})$ is étale with finitely generated group of $K_{s}$-points. By Galois descent, it suffices to show this after extending scalars to some finite Galois extension $L/K$, so we may assume that all of the codimension-$1$ components of $X\backslash U$ are geometrically irreducible. In that case, if the components are $D_{1},\dots,D_{n}$, then the natural map $\psi\colon\oplus_{D}\mathbf{Z}\rightarrow\ker(j^{*})$ is surjective. By Lemma 2.13, $\ker(\psi)$ is a constant group scheme, so $\ker(j^{*})$ is étale (with finitely generated group of $K_{s}$-points). The only if direction of the lemma now follows from Lemma 2.9.
Now we turn to the proof of the if direction (so $K$ need no longer be separably closed). Write $\operatorname{Pic}^{+}_{U/K}\simeq G/E$ with $G$ a smooth commutative $K$-group scheme with finitely generated component group, and $E\subset G^{t}$tale with $E(K_{s})$ free of finite rank, and let $H$ be defined by the following Cartesian square:
@@
Lemma 2.10 implies that $H$ is a (smooth) scheme, hence a smooth commutative $K$-group scheme. Its component group is finitely generated, because the same holds for $G$ and $\ker(j^{*})$. We have $\operatorname{Pic}^{+}_{X/K}\simeq H/E$, and Lemma 2.9 implies that $\operatorname{Pic}^{+}_{X/K}$ is almost-representable. ∎
Next we show that almost-representability descends through finite purely inseparable extensions.
Lemma 5.2.
Let $L/K$ be a finite purely inseparable extension field, let $X$ be a regular $K$-scheme of finite type, and suppose that $Y:=(X_{L})_{\rm{red}}$ is regular. If $\operatorname{Pic}^{+}_{Y/L}$ is almost-representable, then so is $\operatorname{Pic}^{+}_{X/K}$.
Proof.
[…]
@@
If $\operatorname{Pic}^{+}_{X_{L}/L}$ is almost-representable for some finite Galois extension $L/K$, then so is $\operatorname{Pic}^{+}_{X_{L}/K}=\operatorname{R}_{L/K}(\operatorname{Pic}^{+}_{X_{L}/L})$, hence, by Proposition 2.15, so is $\operatorname{Pic}^{+}_{X/K}$. We may therefore assume that each component of $X$ is geometrically irreducible, and therefore, working component by component, that $X$ is geometrically irreducible. Then by [deJ2, Th. 5.9], there is a generically Galois alteration $Y\rightarrow X_{K_{\rm{perf}}}$ such that $\overline{Y}$ admits a smooth compactification. (That theorem applies when ${\rm{dim}}(X)>0$, but of course the 00-dimensional case is trivial.) The data descends from $K_{\rm{perf}}$ down to some finite purely inseparable extension $L/K$. Thus, there exist a finite purely inseparable extension $L/K$ and a proper, dominant, generically Galois morphism of $f\colon Y\rightarrow(X_{L})_{\rm{red}}$ from an integral $L$-scheme $Y$ such that $Y$ admits an $L$-smooth compactification $\overline{Y}$. Shrinking $X$ using lemma 5.1, we may assume that $f$ is finite Galois and that $(X_{L})_{\rm{red}}$ is regular.
The Picard functor $\operatorname{Pic}_{\overline{Y}/L}$ of the proper $L$-scheme $\overline{Y}$ is represented by a locally finite type $L$-scheme, which by the classical theorem of the base has finitely generated component group. Furthermore, this functor is the étale (rather than merely fppf) sheafification of the naive Picard functor $T\mapsto\operatorname{Pic}(\overline{Y}\times_{L}T)$ on the category of $L$-schemes, because $\overline{Y}$ is geometrically reduced. When we restrict to the category of smooth $L$-schemes, $\operatorname{Pic}_{\overline{Y}/L}$ agrees with the functor represented by its maximal smooth $L$-subgroup scheme. It follows that $\operatorname{Pic}^{+}_{\overline{Y}/L}$ is almost-representable (representable even). By Lemma 5.1, so too is $\operatorname{Pic}^{+}_{Y/L}$. Proposition 2.15 then implies that $\operatorname{Pic}^{+}_{(X_{L})_{\rm{red}}/L}$ is almost-representable, and Lemma 5.2 then implies the same for $\operatorname{Pic}^{+}_{X/K}$. ∎
Next we prove the uniqueness of the connected group $G^{0}$ arising in a presentation $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of $\operatorname{Pic}^{+}_{X/K}$ as an almost-representable functor.
@@
We may assume that $n=1$. Because $G$ is normal, ${\rm{H}}^{1}(G,\mathbf{Z})=0$. By [DG, Ch. II, §3, Prop. 2.3], therefore, the commutative extensions of $G$ by $\mathbf{Z}$ may be canonically identified with a subgroup of the second Hochschild cohomology group ${\rm{H}}^{2}_{0}(G,\mathbf{Z})$. Because $G$ is geometrically connected, $G\times G$ is connected, so any $2$-cocycle $G\times G\rightarrow\mathbf{Z}$ is constant, hence is a coboundary. It follows that $\operatorname{Ext}^{1}(G,\mathbf{Z})=0$. ∎
@@
Thanks to Galois descent and the uniqueness aspect of the lemma, we may assume that $E\simeq\mathbf{Z}^{n}$. The obstruction to lifting $\phi$ lies in $\operatorname{Ext}^{1}(G,E)$, which vanishes by Lemma 5.4. Given two lifts $\psi$, $\psi^{\prime}$, their difference yields a homomorphism $G\rightarrow E$, which must be trivial because $G$ is connected and $E$ is étale. ∎
@@
((((Theorem 1.3)))) If $K$ is a field and $X$ is a regular $K$-scheme of finite type, then we have an isomorphism $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$, where $G$ is a smooth commutative $K$-group scheme with finitely generated component group and $E\subset G^{t}$ is an étale $K$-group scheme such $E(K_{s})$ is a free abelian group of finite rank. Furthermore, $G^{0}$ is unique up to unique isomorphism.
@@
Combine Theorem 5.3 and Proposition 5.6. ∎
@@ 6 Structure of the restricted Picard functor
where $f\colon X\times S\rightarrow S$ is the projection, and the cohomology is étale. If $X(S)\neq\emptyset$, then we have an exact sequence
@@
Our next goal is to show that $\operatorname{Pic}^{+}_{X/K}$ has no torus component and contains no copy of ${\mathbf{G}}_{\rm{a}}$ under mild hypotheses on $X$ (for instance, smoothness suffices). To this end, we prove the following proposition.
@@
Let $X$ be a normal $K$-scheme of finite type such that each component of $X_{K_{s}}$ admits a $K_{s}$-point. Then there is no non-constant map $\mathbf{A}^{1}-\{0\}\rightarrow\operatorname{Pic}^{+}_{X/K}$.
@@
We may assume that $K$ is separably closed. Let $f\colon X\rightarrow{\rm{Spec}}(K)$ denote the structure map. Let $S:=\mathbf{A}^{1}-\{0\}$ for notational convenience. Suppose given an element $\alpha\in\operatorname{Pic}^{+}_{X/K}(S)$. By Corollary 6.2, the obstacle $\beta$ to $\alpha$ arising from a line bundle on $X\times S$ lives in ${\rm{H}}^{2}(S,\mathbf{Z}^{n})$ for some $n$. By Lemma 6.3, we have a functorial isomorphism ${\rm{H}}^{2}(S,\mathbf{Z})\simeq{\rm{H}}^{1}(S,\mathbf{Q}/\mathbf{Z})$. The latter agrees with ${\rm{Hom}}_{\rm{cts}}(\pi_{1}(S),\mathbf{Q}/\mathbf{Z})$. Now any finite connected étale cover of $S$ is given by a map $S\rightarrow S$. Passing to such a cover, therefore, we may assume that $\beta=0$, hence $\alpha$ arises from a line bundle on $X\times S$. Now $\operatorname{Pic}(X\times S)=\operatorname{Pic}(X)$ [EGAIV4, Cor. 21.4.11], so $\alpha$ actually arises from an element of $\operatorname{Pic}^{+}_{X/K}(K)$. That is, $\alpha$ is constant. ∎
@@
First suppose that $K$ is perfect. Then we may pass to the separable closure and assume that $K$ is algebraically closed. By Chevalley’s Theorem, $G^{0}$ is an extension of an abelian variety by a smooth connected commutative linear algebraic group $L$. Because $K$ is algebraically closed, $L={\mathbf{G}}_{\rm{m}}^{n}\times\mathbf{A}^{m}$ as a $K$-scheme for some $n,m\geq 0$. Proposition 6.4 therefore implies that $L=0$, so $G^{0}$ is an abelian variety.
Next suppose that $K$ has characteristic $p>0$. By [CGP, Th. A.3.9], one has an exact sequence of smooth connected $K$-group schemes
$0\longrightarrow A\longrightarrow G^{0}\longrightarrow L\longrightarrow 0$
with $A$ semi-abelian and $L$ affine. We claim that $A$ is an abelian variety and that $L$ is unipotent. We may assume that $K$ is separably closed. If $A$ were not an abelian variety, we would obtain a contradiction to Proposition 6.4. If $L$ were not unipotent, then it would contain a copy of ${\mathbf{G}}_{\rm{m}}$, hence we would obtain a subgroup $H$ of $G^{0}$ which is an extension of ${\mathbf{G}}_{\rm{m}}$ by an abelian variety. We claim that any such extension contains a copy of ${\mathbf{G}}_{\rm{m}}$, which would once again contradict Proposition 6.4.
To see the claim, note that $H$ defines an element of $\operatorname{Ext}^{1}({\mathbf{G}}_{\rm{m}},A)$. By Chevalley, over the perfect closure of $K$, $H$ must contain a copy of ${\mathbf{G}}_{\rm{m}}$ which necessarily maps isogenously onto the ${\mathbf{G}}_{\rm{m}}$ quotient. Because any endomorphism of ${\mathbf{G}}_{\rm{m}}$ over an extension field is already defined over $K$, we thus obtain a $K$-isogeny $\phi\colon{\mathbf{G}}_{\rm{m}}\rightarrow{\mathbf{G}}_{\rm{m}}$ such that $H$ splits upon pullback along $\phi$. Then the splitting composed with the map $\phi^{*}(H)\rightarrow H$ provides a nonzero homomorphism ${\mathbf{G}}_{\rm{m}}\rightarrow H$, whose image is then a copy of ${\mathbf{G}}_{\rm{m}}$ sitting inside $H$. ∎
@@
If $X$ is a smooth $K$-scheme of finite type such that each connected component of $X_{K_{s}}$ admits no non-constant units, then there is no non-constant map from a smooth unirational $K$-scheme to $\operatorname{Pic}^{+}_{X/K}$. In particular, there is no such map to $G$, where $G/E$ is a presentation of $\operatorname{Pic}^{+}_{X/K}$ as an almost-representable functor.

### v4

@@ 5 Almost-representability of the restricted Picard functor
In this section we prove the almost-representability of the restricted Picard functor for regular $K$-schemes of finite type.
Lemma 5.1.
Proof.
As a dévissage step, we show that, in proving Theorem 1.3, we are free to restrict to an open subscheme.
@@
Let $j\colon U\rightarrow X$ denote the inclusion, and consider the restriction map $j^{*}\colon\operatorname{Pic}^{+}_{X/K}\rightarrow\operatorname{Pic}^{+}_{U/K}$. For every smooth $K$-scheme $T$, $X\times T$ is regular, hence $\operatorname{Pic}(X\times T)\rightarrow\operatorname{Pic}(U\times T)$ is surjective. It follows that $j^{*}$ is a surjection of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$. We claim that $\ker(j^{*})$ is étale with finitely generated group of $K_{s}$-points, which will pprove the only if direction of the lemma. By Galois descent, it suffices to show this after extending scalars to some finite Galois extension $L/K$, so we may assume that all of the codimension-$1$ components of $X\backslash U$ are geometrically irreducible. In that case, if the components are $D_{1},\dots,D_{n}$, then the natural map $\psi\colon\oplus_{D}\mathbf{Z}\rightarrow\ker(j^{*})$ is surjective. By Lemma 5.1, $\ker(\psi)$ is a constant group scheme, so $\ker(j^{*})$ is étale (with finitely generated group of $K_{s}$-points).
Now we turn to the proof of the if direction (so $K$ need no longer be separably closed). Write $\operatorname{Pic}^{+}_{U/K}\simeq G/E$ with $G$ a smooth commutative $K$-group scheme with finitely generated component group, and $E\subset G$tale with $E(K_{s})$ of finite rank, and let $H$ be defined by the following Cartesian square:
@@
Lemma 2.10 implies that $H$ is a (smooth) scheme, hence a smooth commutative $K$-group scheme. Its component group is finitely generated, because the same holds for $G$ and $\ker(j^{*})$. We thus have that $\operatorname{Pic}^{+}_{X/K}\simeq H/E$ is almost-representable. ∎
@@
Working component by component, we may assume that $X$ is irreducible. Then by [deJ2, Cor. 5.15], there are alterations $Y_{2}\xrightarrow{f_{2}}Y_{1}\xrightarrow{f_{1}}(X_{K_{{\rm{perf}}}})_{\rm{red}}$ with $f_{2}$ generically Galois and $f_{1}$ generically purely inseparable (that is, the function field extension $k(Y_{1})/k((X_{K_{{\rm{perf}}}})_{\rm{red}})$ is purely inseparable) and such that $Y_{2}$ admits a regular, hence $K_{{\rm{perf}}}$-smooth, compactification $\overline{Y}_{2}$. The data descends from $K_{\rm{perf}}$ down to some finite purely inseparable extension $L/K$, and then renaming and postcomposing with the map $(X_{L})_{\rm{red}}\rightarrow X$, we find that there exist a finite purely inseparable extension $L/K$ and dominant generically finite morphisms $Y_{2}\xrightarrow{f_{2}}Y_{1}\xrightarrow{f_{1}}X$ between integral finite type $K$-schemes such that (1) $Y_{2}$ admits an $L$-smooth compactification $\overline{Y}_{2}$; (2) $f_{2}$ is generically Galois; and (3) $f_{1}$ is generically purely inseparable. Shrinking $X$ using lemma 5.2, we may assume that $f_{2}$ is finite Galois and that $f_{1}$ is finite flat radicial.
The Picard functor $\operatorname{Pic}_{\overline{Y}_{2}/L}$ of the proper $L$-scheme $\overline{Y}_{2}$ is represented by a locally finite type $L$-scheme, which by the classical theorem of the base has finitely generated component group. Furthermore, this functor is the étale (rather than merely fppf) sheafification of the naive Picard functor $T\mapsto\operatorname{Pic}(\overline{Y}_{2}\times_{L}T)$ on the category of $L$-schemes, because $\overline{Y}_{2}$ is geometrically reduced. (Work on each geometric component of $\overline{Y}_{2}$, and use [Kle, Th. 9.2.5, Exercise 9.3.11].) When we restrict to the category of smooth $L$-schemes, $\operatorname{Pic}_{\overline{Y}_{2}/L}$ agrees with the functor represented by its maximal smooth $L$-subgroup scheme. It follows that $\operatorname{Pic}^{+}_{\overline{Y}_{2}/L}$ is almost-representable (representable even). By Lemma 5.2, so too is $\operatorname{Pic}^{+}_{Y_{2}/L}$. Proposition 2.14 then implies that $\operatorname{Pic}^{+}_{Y_{1}/L}$ is almost-representable. Now, just by definition,
$(\operatorname{Pic}^{+}_{Y_{1}/K})^{\rm{naive}}=\operatorname{R}_{L/K}((\operatorname{Pic}^{+}_{Y_{1}/L})^{\rm{naive}}).$
Because $L/K$ is finite radicial surjective, it induces an equivalence of étale sites [SGA, Exp. IX, Th. 4.10], hence $\operatorname{R}_{L/K}$ commutes with étale sheafification, so we obtain an isomorphism
$\operatorname{Pic}^{+}_{Y_{1}/K}\simeq\operatorname{R}_{L/K}(\operatorname{Pic}^{+}_{Y_{1}/L}).$
Since pushforward through the finite morphism ${\rm{Spec}}(L)\rightarrow{\rm{Spec}}(K)$ is an exact functor between étale sites, it preserves almost-representability, so $\operatorname{Pic}^{+}_{Y_{1}/K}$ is almost-representable. Therefore, combining Propositions 3.4, 4.1, and 4.4, we obtain that $\operatorname{Pic}^{+}_{X/K}$ is almost-representable. ∎
Next we will prove the uniqueness of the connected group $G^{0}$ arising in a presentation $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of $\operatorname{Pic}^{+}_{X/K}$ as an almost-representable functor. We require a couple of lemmas.
@@
We may assume that $n=1$. Because $G$ is normal, ${\rm{H}}^{1}(G,\mathbf{Z})=0$. By [DG, Ch. II, §3, Prop. 2.3] or [Con1, Prop. B.2.5], therefore, the commutative extensions of $G$ by $\mathbf{Z}$ may be canonically identified with a subgroup of the second Hochschild cohomology group ${\rm{H}}^{2}_{0}(G,\mathbf{Z})$. Because $G$ is geometrically connected, $G\times G$ is connected, so any $2$-cocycle $G\times G\rightarrow\mathbf{Z}$ is constant, hence is a coboundary. It follows that $\operatorname{Ext}^{1}(G,\mathbf{Z})=0$. ∎
@@
We may replace $G$ by $G/\mathscr{D}G$ and thereby assume that $G$ is commutative. Thanks to Galois descent and the uniqueness aspect of the lemma, we may assume that $E\simeq\mathbf{Z}^{n}$. The obstruction to lifting $\phi$ lies in $\operatorname{Ext}^{1}(G,E)$, which vanishes by Lemma 5.4. Given two lifts $\psi$, $\psi^{\prime}$, their difference yields a homomorphism $G\rightarrow E$, which must be trivial because $G$ is connected and $E$ is étale. ∎
@@
((((Theorem 1.3)))) If $K$ is a field and $X$ is a regular $K$-scheme of finite type, then we have an isomorphism $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$, where $G$ is a smooth commutative $K$-group scheme with finitely generated component group and $E\subset G^{t}$ is an étale $K$-group scheme such that $E(K_{s})$ is a free abelian group of finite rank. Furthermore, $G^{0}$ is unique up to unique isomorphism.
@@
Combine Theorem 5.3, Lemma 2.9, and Proposition 5.6. ∎
@@ 6 Structure of the restricted Picard functor
where $f\colon X\times S\rightarrow S$ is the projection, and the cohomology is étale. If $X(S)\neq\emptyset$, then the above sequence induces an exact sequence
@@
The following proposition says that a map from a connected group into $\operatorname{Pic}^{+}_{X/K}$ arises from a line bundle after passing to an isogenous cover.
@@
Let $K$ be a separably closed field, and let $X$ be a normal $K$-scheme of finite type each component of which admits a $K$-point. Then for any smooth connected $K$-group scheme $G$, and any homomorphism $\phi\colon G\rightarrow\operatorname{Pic}^{+}_{X/K}$ of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$, there is a finite étale $K$-homomorphism $\psi\colon H\rightarrow G$ from a smooth connected $K$-group scheme $H$ such that $\phi\circ\psi$ arises from a line bundle on $X\times_{K}H$.
@@
Combining Corollary 6.2 and Lemma 6.3, the obstacle to a map $g\colon S\rightarrow\operatorname{Pic}^{+}_{X/K}$ from a finite type smooth $S/K$ arising from a line bundle on $X\times_{K}S$ is an element in $\Delta_{g}\in{\rm{H}}^{1}(S,\mathbf{Z}^{n})$ for some $n\geq 0$ (independent of $S$) which is associated in a functorial manner to $g$. Furthermore, this obstacle is additive in $g$: Given $g_{i}\colon S\rightarrow\operatorname{Pic}^{+}_{X/K}$, $i=1,2$, the obstacle for $g_{1}+g_{2}\colon S\rightarrow\operatorname{Pic}^{+}_{X/K}$ is the sum of the obstacles for the $g_{i}$.
We apply the above with $S=G\times G$, $g_{i}=\phi\circ\pi_{i}$ with $\pi_{i}\colon G^{2}\rightarrow G$ the projections. The map $g_{1}+g_{2}$ coincides with $\phi\circ m$, where $m\colon G^{2}\rightarrow G$ is addition. Thus we obtain
$m^{*}\Delta_{\phi}=\Delta_{\phi\circ m}=\Delta_{g_{1}+g_{2}}=\Delta_{g_{1}}+\Delta_{g_{2}}=\pi_{1}^{*}\Delta_{\phi}+\pi_{2}^{*}\Delta_{\phi}.$
That is,
$m^{*}\Delta_{\phi}=\pi_{1}^{*}\Delta_{\phi}+\pi_{2}^{*}\Delta_{\phi}.$ (6.1)
Now $\Delta_{\phi}\in{\rm{H}}^{1}(G,A)\subset{\rm{H}}^{1}(G,\mathbf{Q}/\mathbf{Z})$ for some finite constant subgroup scheme $A\subset\mathbf{Q}/\mathbf{Z}$.
Let $h\colon T\rightarrow G$ be the $A$-torsor corresponding to $\Delta_{\phi}$. Note that pullback along $h$ kills $\Delta_{\phi}$, so the map $\phi\circ h\colon T\rightarrow\operatorname{Pic}^{+}_{X/K}$ arises from a line bundle on $X\times_{K}T$. Because $K$ is separably closed, ${\rm{H}}^{1}(K,A)=0$, so there is a point $e\in T(K)$ lying over $1\in G(K)$. We claim that equation (6.1) implies that $T$ admits a $K$-group scheme structure with identity $e$ making $h$ into a homomorphism with kernel $A$. Indeed, this follows from the same argument as in the proof of [C-T, Th. 4.12]. That result treats the case when $A$ is of multiplicative type instead of constant, and uses Rosenlicht’s theorem that then any map $G^{3}\rightarrow A$ sending identity to identity is a character. In fact, all that the argument really uses is that any map $G^{3}\rightarrow S$ is a sum $\sum_{i=1}^{3}\gamma_{i}\circ\pi_{i}$, with $\pi_{i}\colon G^{3}\rightarrow G$ the projection and some $\gamma_{i}\colon G\rightarrow A$. This holds in our setting with $G$ connected and $A$tale because in fact any map $G^{3}\rightarrow A$ must then be constant.
So $T$ is a group scheme admitting a finite étale $K$-homomorphism to $G$. It follows that $T^{0}\rightarrow G$ is also finite étale. Since $T^{0}\rightarrow G$ naturally factors through the inclusion $T^{0}\rightarrow T$, the composition $T^{0}\rightarrow G\rightarrow\operatorname{Pic}^{+}_{X/K}$ is induced by a line bundle on $X\times_{K}T^{0}$. ∎
The following result says that, quite generally, $\operatorname{Pic}^{+}_{X/K}$ does not admit nontrivial maps from tori.
Proposition 6.5.
Let $X$ be a normal $K$-scheme of finite type such that each component of $X_{K_{s}}$ admits a $K_{s}$-point. Then there is no nonzero homomorphism $T\rightarrow\operatorname{Pic}^{+}_{X/K}$ with $T$ a $K$-torus.
Proof.
We may extend scalars to $K_{s}$ and may therefore assume that $T={\mathbf{G}}_{\rm{m}}$. Let $f\colon{\mathbf{G}}_{\rm{m}}\rightarrow\operatorname{Pic}^{+}_{X/K}$ be a homomorphism. By Proposition 6.4, there is an isogeny $\psi\colon H\rightarrow{\mathbf{G}}_{\rm{m}}$ with $H$ smooth and connected such that $f\circ\psi$ arises from a line bundle. The $K$-group $H$ must also be isomorphic to ${\mathbf{G}}_{\rm{m}}$, hence we are reduced to the case in which $f$ arises from a line bundle. But pullback along the projection induces an isomorphism $\operatorname{Pic}(X)\xrightarrow{\sim}\operatorname{Pic}(X\times_{K}{\mathbf{G}}_{\rm{m}})$ [EGA, IV4, Cor. 21.4.11], so $f$ is a constant map, hence 00. ∎
Proposition 6.6.
If $K$ is a perfect field, and $X$ is a normal $K$-scheme of finite type, then there is no nonzero homomorphism ${\mathbf{G}}_{\rm{a}}\rightarrow\operatorname{Pic}^{+}_{X/K}$.
Proof.
We may extend scalars and assume that $K=\overline{K}$. Let $f\colon{\mathbf{G}}_{\rm{a}}\rightarrow\operatorname{Pic}^{+}_{X/K}$ be a homomorphism. By Proposition 6.4, there is an isogeny $\psi\colon H\rightarrow{\mathbf{G}}_{\rm{a}}$ with $H$ smooth and connected such that $f\circ\psi$ arises from a line bundle. Because $K$ is algebraically closed, $H\simeq{\mathbf{G}}_{\rm{a}}$. Thus we are reduced to the case in which $f$ arises from a line bundle. But pullback induces an isomorphism $\operatorname{Pic}(X)\xrightarrow{\sim}\operatorname{Pic}(X\times_{K}{\mathbf{G}}_{\rm{a}})$ [EGA, IV4, Cor. 21.4.11], so $f$ is a constant map, hence 00. ∎
Remark 6.7.
Proposition 6.6 fails in general even for smooth $X$ without the perfectness assumption on $K$. More precisely, for every imperfect field $K$, there is a smooth finite type $K$-scheme $X$ such that $\operatorname{Pic}^{+}_{X/K}$ is (represented by) a nontrivial split unipotent $K$-group scheme; see Example 6.14.
@@
First suppose that $K$ is perfect. By Chevalley’s theorem [Con2], we may write $G^{0}$ as an extension
$0\longrightarrow L\longrightarrow G^{0}\longrightarrow A\longrightarrow 0$
with $L/K$ a connected linear algebraic group and $A/K$ an abelian variety. We must show that $L=0$. Because $L$ contains no nontrivial torus by Proposition 6.5, $L=U$ is unipotent [Bor, IV, Cor. 11.5(2)]. Because $K$ is perfect, every smooth connected unipotent $K$-group is split [CGP, Cor. B.2.7, B.3.3], so we are done by Proposition 6.4.
Now consider the imperfect case. By [CGP, Th. A.3.9] and Proposition 6.5, we have an exact sequence
$1\longrightarrow A\longrightarrow G^{0}\longrightarrow L\longrightarrow 1$
with $A$ an abelian variety and $L$ linear algebraic. We must show that $L$ does not contain a nontrivial torus (and is therefore unipotent). This follows from the fact that $G^{0}$ does not contain a nontrivial torus and [CGP, Prop. C.4.5(2)]. ∎
We now prove Theorem 1.4.
Theorem 6.9.
((((Theorem 1.4)))) Let $K$ be a field and $X$ a regular $K$-scheme of finite type with dense smooth locus. Assume that any map from $X$ into an abelian variety $A$ factors through a finite subscheme of $A$. Then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative $K$-group scheme with finitely generated component group. Furthermore, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is unipotent.
Proof.
Let $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ be an almost-representable presentation of $\operatorname{Pic}^{+}_{X/K}$ (Theorem 5.3) as in Lemma 2.9. By Corollary 6.8, $G^{0}$ is an extension of a smooth unipotent group scheme by an abelian variety $A$. The key point is to show that $A=0$. We claim that we may assume that $K$ is separably closed. The only unclear point is that for $X$ not to admit a map with infinite image into an abelian variety is preserved upon passage to $K_{s}$. For this, we simply note that, given a finite separable extension $L/K$, and a map $X_{L}\rightarrow B$ (with $B$ an abelian variety over $L$) with infinite image, the corresponding map $X\rightarrow\operatorname{R}_{L/K}(B)$ also has infinite image, where $\operatorname{R}_{L/K}$ denotes Weil restriction. Since $\operatorname{R}_{L/K}(B)$ is also an abelian variety, we may indeed assume that $K$ is separably closed.
We may also assume that $X$ is nonempty and connected, hence irreducible. We will show that there is no nonzero homomorphism $B\rightarrow G^{0}$ with $B$ an abelian variety, hence $A=0$. Thanks to Proposition 6.4, we may by passing to an isogenous cover of $B$ assume that the homomorphism $B\rightarrow G/E\simeq\operatorname{Pic}^{+}_{X/K}$ comes from a line bundle $\mathscr{L}$ on $X\times_{K}B$. Choose a point $x_{0}\in X(K)$ (which exists because $X$ has nonempty smooth locus and $K$ is separably closed). We are free to modify $\mathscr{L}$ by a line bundle pulled back from $B$ and thereby assume that $\mathscr{L}|_{x_{0}}\simeq\mathcal{O}_{B}$.
Now $\mathscr{L}$ defines a map $X\rightarrow\operatorname{Pic}_{B/K}$ which sends $x_{0}$ to 00. Since $X$ is connected, therefore, it defines a map $X\rightarrow\widehat{B}$ to the dual abelian variety. By hypothesis, this map has finite image. Because $X$ is integral and $X(K)\neq\emptyset$, it is a constant map to a point $b\in B(K)$. This means that $\mathscr{L}\in\operatorname{Pic}(X)\oplus\operatorname{Pic}(B)$, so the map $B\rightarrow\operatorname{Pic}^{+}_{X/K}$ is constant, hence 00. This in turn means that the map $h\colon B\rightarrow G$ factors through $E\subset G$. Because $B$ is connected and $E$tale, it follows that $h=0$, so we conclude that $A=0$. By Corollary 6.8, it follows that $G^{0}$ is smooth and unipotent, and even trivial if $K$ is perfect. Thus $G^{t}$ is torsion. It follows that $E\subset G^{t}$ is trivial. Thus $\operatorname{Pic}^{+}_{X/K}\simeq G$ is a smooth $K$-group scheme with unipotent identity component. ∎
The situation is simplified when $X$ has no nonconstant units.
@@
If $X$ is a regular $K$-scheme of finite type with dense smooth locus satisfying the property that each connected component of $X_{K_{s}}$ admits no non-constant units, then there is no non-constant map from a smooth unirational $K$-scheme to $\operatorname{Pic}^{+}_{X/K}$. In particular, there is no such map to $G$, where $G/E$ is a presentation of $\operatorname{Pic}^{+}_{X/K}$ as an almost-representable functor.